1. If 2.0x10-4moles of S2O82-in 50 mL of solution is consumed in 188 seconds, what is the rate of consumption of S2O82-?
Answer #1

Rate = AC - mol/lit. RateNelocity of chemical reaction : The rate of change of concentration with time of different chemical|2x10-4 moly of ₂ og 2- Il volume = 50mL = 50x10-32 time = 108 seg Concentration of S₂ 92-= moles Volumelini) 2x 10-4 sox 10-